In the small town of Geneva, there are five firms that make watches. The firm's respective output levels are (watches per year):

first firm 30,

second firm 20,

third firm 20

fourth firm 20

fifth firm 10.

Given the information above, compute the four-firm concentration ratio and Herfindahl index for the watch industry in Geneva. Based on your calculations, is the watch industry in Geneva monopolistically competitive or is it oligopolistic? Explain your answer.
1
Expert's answer
2020-04-02T09:39:29-0400

The four-firm concentration ratio is calculated by adding the market shares of the four largest firms: in this case, 30 + 20 + 20 + 20 = 90. This concentration ratio would can be considered especially high, because the largest four firms have 90% of the market. Herfindahl index is calculated by squaring the market share of each firm competing in a market and then summing the resulting numbers. In this case market share of the 1st firm is 30% (30/(30+20+20+20+10)*100%=30%), 2nd, 3rd, 4th - 20%, 5th - 10%. Herfindahl index is

"30*30+(20*20)*3+10*10=2,200"

Based on this calculations, the watch industry in Geneva is a moderately concentrated marketplace, (less than 2,500), but taking into account the four-firm concentration ratio the market is oligopolistic.